OpenStudy (anonymous):

lim as (x->0) of (tan(2 x))/(tan(pi x)) = 2/pi can some please explain the process of arriving at this limit without using the l'hospital rule

OpenStudy (anonymous):

l'hospital's rule is VERY important in calculus... and i WOULD use it for this problem...

OpenStudy (anonymous):

we are not allowed to use it thus far in calculus because we havent been taught it. Im in calc one second week. i cant find away around it though

OpenStudy (anonymous):

You can use the fact that tan(c*x) is approximated by the line c*x when x is close to zero

OpenStudy (anonymous):

@zdstrong you have to first rewrite the numerator and denominator into terms of sin and cos

OpenStudy (anonymous):

you can do it shi's way also... longer however:)

OpenStudy (anonymous):

so \[\frac{ \tan(2x) }{\tan(\pi x) } = \frac{ \frac{ \sin(2x) }{ \cos(2x) } }{ \frac{ \sin(\pi x) }{ \cos(\pi x) } } = \frac{ \frac{ \sin(2x) }{ 1 }*\frac{ 1 }{ \cos(2x) } }{ \frac{ \sin(\pi x) }{ 1 }*\frac{ 1 }{ \cos(\pi x) } } \] still trying to solve the problem?

OpenStudy (anonymous):

@Shi you need to use sinc(x) ... not sure if you were going that way eventually...

OpenStudy (anonymous):

haha I don't think im aware of your method Algebraic, though this method is probably how they want it solved

OpenStudy (anonymous):

no.. afterward I just multiply the numerator by 2x/2x and the denominator by (pi x)/(pi x)

OpenStudy (anonymous):

\[\sin (2*x) / \sin (\pi*x) * (2*x/2*x) / (\pi*x/\pi*x)\]

OpenStudy (anonymous):

what is sinc(x)?? I dont think ive used it before and i want to learn haha

OpenStudy (anonymous):

yes:) we're doing the same thing... the tangent thing is easier though... linearization of tan(c*x) = c*sec^2(c*x) for x=0

OpenStudy (anonymous):

sinc(x) is sin(x) / x

OpenStudy (anonymous):

and the limit is 1 as x->0

OpenStudy (anonymous):

same thing you were doing I think...

OpenStudy (anonymous):

ah, don't think I knew that, and not sure if that was taught at the level of this problem

OpenStudy (anonymous):

it's the only other way to do it afaik... what were you going to do?

OpenStudy (anonymous):

I was talking about the linearization of tan haha ok just to clear everything up for op the last step looks like